OpenStudy (anonymous):

Write the expression in standard form. five divided by quantity three minus fifteen i.

OpenStudy (anonymous):

\[\frac{ 5 }{ 3-15i }\]

OpenStudy (anonymous):

\[\frac{ 5 }{ 3}+\frac{ 5 }{ 15i}?\]

OpenStudy (anonymous):

\[\frac{ 5 }{ 3}+\frac{ 1 }{ 3i }\]

OpenStudy (anonymous):

Not an answer choice

OpenStudy (anonymous):

\[\frac{5 (3+15 i)}{(3-15 i) (3+15 i)}=\frac{1}{234} (15+75 i)=\frac{1}{78} (5+25 i) \]